OpenStudy (anonymous):

Use the distributive property to multiply. 3(3x+4y+2z)= ?

OpenStudy (anonymous):

multiply each term by 3

OpenStudy (uri):

I was about to type that,lol ^

OpenStudy (anonymous):

what is \(3\times 3x\) ?

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

k good how about \(3\times 4y\)?

OpenStudy (anonymous):

12y

OpenStudy (anonymous):

one more \[3\times 2z\]

OpenStudy (anonymous):

6z

OpenStudy (anonymous):

put it all together and what do you get?

OpenStudy (anonymous):

9x12y6z

OpenStudy (anonymous):

think you are missing some plus signs

OpenStudy (anonymous):

9x+12y+6z?

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

don't worry, the questions get harder

OpenStudy (anonymous):

That worries me @satellite73